# Data Analogies Enable Efficient Cross-Embodiment Transfer

### 06 Mar 2026 v1

@@ Abstract
Generalist robot policies are trained on demonstrations collected across a wide variety of robots, scenes, and viewpoints. Yet it remains unclear how to best organize and scale such heterogeneous data so that it genuinely improves performance in a given target setting. In this work, we ask: what form of demonstration data is most useful for enabling transfer across robot set-ups? We conduct controlled experiments that vary end-effector morphology, robot platform appearance, and camera perspective, and compare the effects of simply scaling the number of demonstrations against systematically broadening the diversity in different ways. Our simulated experiments show that while perceptual shifts such as viewpoint benefit most from broad diversity, morphology shifts benefit far less from unstructured diversity and instead see the largest gains from data analogies, i.e. paired demonstrations that align scenes, tasks, and/or trajectories across different embodiments. Informed by the simulation results, we improve real-world cross-embodiment transfer success by an average of $22.5\%$ over large-scale, unpaired datasets by changing only the composition of the data.
@@ I Introduction
We study how data collection strategies can affect cross-embodiment transfer, when the target robot has only a small amount of collected data. Our investigation considers basic imitation learning on cross-embodiment data, without any special architectural or algorithmic changes such as in-painting. While data diversity is known to contribute to generalization, we also consider an orthogonal aspect of the data, which is the degree the data contains analogous data across robot set-ups. These data analogies entail demonstrations from different robot embodiments with the similar environments, tasks, and/or trajectories, and have the potential to more explicitly teach the model how different robots relate to each other. See Figure 1 for an example.
@@ II Related Works
While these model-centric approaches to alignment and representation have shown promise, they largely treat the underlying data distribution as a given. This overlooks a crucial, emerging direction: focusing on the data distribution itself. For example, recent works have found that carefully collecting and rebalancing data distributions for training VLAs can unlock gains in generalization and transfer [15, 16, 37, 33, 17]. We argue that this data-centric principle is particularly crucial for cross-embodiment learning. Prior works have seen some success in improving robot generalization by expanding the data coverage or diversity of demonstration data [20, 25, 4, 41]. However, it is still unclear what method for scaling are useful to cross the embodiment gap. Analogous to how VLAs provided a high-level glue for generalization, a new data-centric glue is needed to bridge the gaps in embodiment. This work investigates how to scale embodiments by prioritizing the right axes of data coverage and data distribution for high-fidelity cross-embodiment transfer.
@@ VI Limitations
While our results show consistent gains from structured coverage and trajectory pairing, our study has several limitations. First, conclusions are drawn using a $\pi_{0.5}$-style VLA under a fixed few-shot budget; other architectures, larger budgets, or alternative training recipes could shift absolute performance and sometimes the relative gaps. Second, dataset shift remains: OXE and our scenes differ from simulation distributions and from other labs; our compositional strategy may require retuning to new object distributions, sensors, and embodiments. Finally, while our study drew conclusions from diverse simulations, the scope of experiments were within two buildings for the real world. This means that it is unclear to what extent the translation dataset should scale to provide consistent general transfer between two real robots.

### 20 Mar 2026 v2

@@ Abstract
Generalist robot policies are trained on demonstrations collected across a wide variety of robots, scenes, and viewpoints. Yet it remains unclear how to best organize and scale such heterogeneous data so that it genuinely improves performance in a given target setting. In this work, we ask: what form of demonstration data is most useful for enabling transfer across robot set-ups? We conduct controlled experiments that vary end-effector morphology, robot platform appearance, and camera perspective, and compare the effects of simply scaling the number of demonstrations against systematically broadening the diversity in different ways. Our simulated experiments show that while perceptual shifts such as viewpoint benefit most from broad diversity, morphology shifts benefit far less from unstructured diversity and instead see the largest gains from data analogies, i.e. paired demonstrations that align scenes, tasks, and/or trajectories across different embodiments. Informed by the simulation results, we improve real-world cross-embodiment transfer success by an average of $22.5\%$ over large-scale, unpaired datasets by changing only the composition of the data. For further information, please visit https://data-analogies.github.io/.
@@ I Introduction
We study how data collection strategies can affect cross-embodiment transfer, when the target robot has only a small amount of collected data. Our investigation considers basic imitation learning on cross-embodiment data, without any special architectural or algorithmic changes such as in-painting. While data diversity is known to contribute to generalization, we also consider an orthogonal aspect of the data, which is the extent to which the data contains analogous examples across robot set-ups. These data analogies entail demonstrations from different robot embodiments with similar environments, tasks, and/or trajectories, and have the potential to more explicitly teach the model how different robots relate to each other. See Figure 1 for an example.
@@ II Related Works
While these model-centric approaches to alignment and representation have shown promise, they largely treat the underlying data distribution as a given. This overlooks a crucial, emerging direction: focusing on the data distribution itself. For example, recent works have found that carefully collecting and rebalancing data distributions for training VLAs can unlock gains in generalization and transfer [15, 16, 37, 33, 17]. We argue that this data-centric principle is particularly crucial for cross-embodiment learning. Prior works have seen some success in improving robot generalization by expanding the data coverage or diversity of demonstration data [20, 25, 4, 41]. However, it is still unclear what methods for scaling are useful to cross the embodiment gap. Analogous to how VLAs provided a high-level glue for generalization, a new data-centric glue is needed to bridge the gaps in embodiment. This work investigates how to scale embodiments by prioritizing the right axes of data coverage and data distribution for high-fidelity cross-embodiment transfer.
@@ VI Limitations
While our results show consistent gains from structured coverage and trajectory pairing, our study has several limitations. First, conclusions are drawn using a $\pi_{0.5}$-style VLA under a fixed few-shot budget; other architectures, larger budgets, or alternative training recipes could shift absolute performance and sometimes the relative gaps. Second, dataset shift remains: OXE and our scenes differ from simulation distributions and from other labs; our compositional strategy may require retuning to new object distributions, sensors, and embodiments. Finally, while our study drew conclusions from diverse simulations, our real-world experiments were conducted within two buildings. This means that it is unclear to what extent the translation dataset should scale to provide consistent general transfer between two real robots.